calculate the cash conversion cucle and all of its components for Walmart and target for the fiscal year 2014
Answer #1

solution :

1. Cash Conversion Cycle = [(Inventory / Cost of sales) + (Account Receivables / Net Sales) + (Accounts Payable / Cost of Sales)] x 365

For Walmart:
=> [(45,141/355,913) + (6,778/483,521) + (38,410/355,913)] x 365 = 90.80 Days

For Target:
=> [(8,790/51,278) + (1,122/72,618) + (7,759/51,278)] x 365 = 123.44 Days

Account receivable turnover = Net Credit Sales / Average Account Receivables
For Walmart:
=> 483,521/[(6,778+6,677)/2] = 71.87 times

For Target:
=> 72,618 / [(1,122+1,347)/2] = 58.82 times

Account receivable days = 365 / Account receivable turnover
For Walmart:
=> 365 / 71.87 = 5.08 days

For Target:
=> 365 / 58.82 = 6.21 days

Inventory turnover ratio = Cost of goods sold / Average inventory
For Walmart:
=> 355,913 / [(45,141 + 44,858)/2] = 7.91 times

For Target:
=> 72,618 / [(8,790 + 8,766)/2] = 8.27 times

Inventory turnover period = 365 / Inventory turnover ratio
For Walmart:
=> 365 / 7.91 = 46.15 days

For Target:
=> 365 / 8.27 = 44.12 days

Accounts payable turnover ratio = (Cost of goods sold + Ending inventory – Beginning inventory) / Average Accounts payable
For Walmart:
=> [(355,913+45,141-44,858)] / [(38,410+37,415)/2] = 9.40 times

For Target:
=> [(51,278+8,790-8,766)] / [(7,759+7,683)/2] = 6.64 times

Accounts payable turnover period = 365 / Accounts payable turnover
For Walmart:
=> 365 / 9.40 = 38.85 days

For Target:
=> 365 / 6.64 = 54.93 days

any doubts comment me below

thank you